The admission fee at an amusement park is 1.5 dollars for children and 4 dollars for adults. On a certain day, 278 people entered the park, and the admission fees collected totaled 792 dollars. How many children and how many adults were admitted?
A man invests his savings in two accounts, one paying 6 percent and the other paying 10 percent simple interest per year. He puts twice as much in the lower-yielding account because it is less risky. His annual interest is 5500 dollars. How much did he invest at each rate?
Your answer is
Country Day's scholarship fund receives a gift of $ 130000. The money is invested in stocks, bonds, and CDs. CDs pay 3.25 % interest, bonds pay 3.6 % interest, and stocks pay 11.9 % interest. Country day invests $ 15000 more in bonds than in CDs. If the annual income form the investments is $ 7445 , how much was invested in each vehicle?



Answer by baseball35(2)

This is the answer to the amusement park problem! Below are two equations and I defined the variables that are in those two equations.
(c + d = 278)-4
1.5c + 4d = 792
-4c + -4d = -1112
You are going to wind up with an equation like this: -2.5c = -320
C = Children at Amusement park
A = Adults at Amusement park
C = 128
A = 150
Thank you for submitting this problem. I love helping
Thanks